# Time-Series Foundation Model for Value-at-Risk

### v2

@@ 4 Evaluation metrics
where the expected violations are calculated as $E=T\times\alpha$, with $T$ being the total number of observations and $\alpha$ the confidence level of the VaR model. Clearly, an AE ratio of 1 is desirable as it indicates that the model is accurately calibrated. On other hand, a value $>1$ ($<1$) suggests the underestimation (overestimation) of the risk, both of which are problematic in the sense that an underestimate leads to unexpected negative returns, while the latter leads to high capital requirements, thus, both implying capital losses. Alternatively, a model is deemed good if $|1-\text{AE}|$ is close to zero.
An alternative to AE ratio is the proportion of failure test (POF), also known as the unconditional coverage test (UC). which statistically evaluates whether the actual number of VaR violations matches the expected number of violations over time, based on the chosen confidence level. For instance, it calculates the failure rate defined as proportion of number of violations ($A$) to the total number of observation $T$ and examines it is consistence with the models confidence level $\alpha$. More specifically, let $I_{t}$ denotes the indicator function that the realized return was less than the VaR estimate, i.e.,
@@
In the evaluation of VaR forecasts, the quantile score is a commonly used metric and is now standard in the VaR literature (Ehm et al., 2016; Gneiting, 2011; Taylor, 2019). The above mentioned statistical tests are good to evaluate the statistical adequacy of VaR estimates, that is, a model for VaR forecasting is deemed to be an adequate model if we fail to reject the null hypothesis. However, these tests cannot conclude whether an ‘adequate’ model outperforms another ‘adequate’ model. Scoring functions allow us to do an honest assessment of VaR as a perfect VaR estimate would always minimize the scoring function. The quantile score $QS_{t}$ can be defined as follows:
@@
For every instant $t$, a model’s score increases either by (a) $\alpha-1$ fraction of the distance between the realized return and the VaR forecast, if realized return is smaller than the VaR forecast for the same period, or (b) by $\alpha$ fraction of the distance otherwise. This scoring rule is consistent for quantile estimation, meaning that it achieves its minimum value when the quantile forecast $q_{t}$ correctly estimates the true quantile of the distribution. Therefore, models that produce lower average quantile scores in the out-of-sample period are considered superior. Finally, to compare the performance of two models, we use Diebold–Mariano (DM) test (Diebold and Mariano, 2002) with
@@ 5.1 Actual over Expected Ratio
Table 2 reports the summary statistics, across 92 stocks, of the $|1-\text{AE}|$ values for the VaR forecasts at different levels, and the corresponding models in the out-samplex‘ period. Specifically, we report the minimum, mean, median, maximum, standard deviation across all the stocks. Additionally, we report the count of stocks for which each of the considered model was the best (achieved lowest value of $|1-AE|$) or was within top two models (1st-2nd best). Note that, as mentioned before, the VaR forecasts for TimesFM pretrained model are only available at $10\%$, there we have three extra models in the last block of Table 2.
@@
where $i$ denotes the model in the selected row, whereas $j$ denotes the model in the selected column. Further, the formatting in the Tables are as follows: *, ** and *** denotes whether the $t$-test of equal predictive accuracy is rejected at the 5%, 2.5% and 1% level of significance. Table 3 reports the results to $t$-tests. We can observe that the fine-tuned models, in particular, FT1 and FT21, performs remarkably well across all quantile levels, depicted by failure to reject the null hypothesis in the first and second rows of each block. Further, we observe that FT1 outperforms GAS model in forecasting VaR(1%), VaR(5%) and VaR(10%) at 97.5%, 95% and 99% level of significance respectively, and performs atleast as good as GAS model for forecasting Var(2.5%) (refer rows corresponding to GAS model). In addition, we observe that selecting short horizon as prediction length performs better in predicting VaR. What’s more,, the performance of pre-trained models is not so exciting compared to their fine-tuned counterpart. These findings show the potential of time-series foundation models in forecasting value at risk.
@@
Table 2: Summary statistics of the $|1-AE|$ values over the out-of-sample period from January 2015 to September 2023 for the eleven models. Additionally, we report the count of stocks for which each of the considered model was the best (achieved lowest value of $|1-AE|$) or was within top two models (1st-2nd best). In case of a tie, equal ranks were given. The values are highlighted using bold for the best values and italicized for the second-best in each column.
@@
Tables 4, 5 and 6 present the results of these tests, namely UC, CC and DQ respectively, for various confidence levels. The results are also depicted via bar plots in Figures A.1-A.3. For each $VaR_{\alpha}$, we present for how many assets we failed to reject the null hypothesis of correct model performance at various confidence levels. Higher the count better the model. From these Tables, it is evident that the TimesFM shows promising results, in particular, the fine-tuned TimesFM model with one prediction length (FT1) either comfortably beats or perform fairly in comparison to the benchmark models across all the three tests. More specifically, for the UC test, the fine-tuned TimesFM models with various prediction lengths and the Historical model showed the highest number of assets where the null hypothesis was not rejected, indicating that these models had the most consistent unconditional coverage. What’s more remarkable is that the simple rolling method (historical) outperforms the GAS model in UC test. Again, in conditional coverage tests, the FT model with 1 prediction length performed significantly better than all other models. Finally, the performance of FT1 in DQ test is comparable to the benchmark models if not better. In conclusion, the FT1 model demonstrates strong performance, providing robust competition to the GAS model across multiple tests, while TimesFM models with longer prediction lenghts show significant weaknesses, particularly in capturing the conditional and dynamic aspects of the VaR process.
@@ 5.2 Quantile Loss
Table 7 presents the summary statistics of quantile loss scores across 92 stocks. We observe that, at 1% VaR, the mean and median quantile score is consistent across the models, with FT1 and FT21 exhibiting lower mean values, 0.062 and 0.065, respectively. This indicates a better performance of fine-tuned TimesFM models compared to GARCH models (G-EDF, G-t, G-N) that have higher mean values and higher values for maximum quantile score. Further, higher value of standard deviation for GARCH models show that these models have high variability in their VaR forecasts which, in turn, reflects that these models are less stable in extreme market scenarios. These observations remain valid at 2.5% VaR forecasts, with FT1 and FT21 showing lower mean and median quantile scores, while GARCH models still show higher values for maximum quantile score and the standard deviation. For VaR at 5% and 10% levels, the historical model performs similar to fine-tuned TimesFM model whereas GARCH models continue to show higher maximum quantile score and greater variability, thus making them less robust and more sensitive to extreme market conditions. In addition, GAS models obain lower values of mean and median quantile scores, similar to that obtained by FT models across all quantile levels. FT models are thus competitive to the benchmark models, in particular, GAS model which show slightly higher values of maximum quantile loss and higher standard deviation particularly at VaR 5% and 10% levels.
@@
As the quantile scores are not easy to interpret, we calculate skill scores defined as the ratio of each model’s score to the score of a benchmark model for an individual asset, and then, we take the mean across assets to summarise performance of models across multiple assets. More specifically, we report in Table 8, the out-of-sample total quantile loss of each model relative to the benchmark model in the selected row. For instance, for VaR (1%), the first row contains the cross-sectional average of the ratio of out-of-sample total quantile loss of every model relative to the out-of-sample total quantile loss of model FT21. In addition to comparison of the total quantile loss, we also perform Diebold–Mariano test for each pair of model where we test
@@
where $i$ denotes the model in the selected row, whereas $j$ denotes the model in the selected column. Further, the formatting in the Tables are as follows: *, ** and *** denotes whether the Diebold–Mariano (DM) test of equal predictive accuracy is rejected for more than 50% of the assets at the 5%, 2.5% and 1% level of significance.
VaR (1%) VaR (2.5%)
FT1 0.000 1.061 1.123 1.097 1.131 1.095 1.005 1.109 0.000 1.045 1.090 1.138 1.142 1.141 0.998 1.078
FT21 0.944 0.000 1.058 1.034 1.066 1.032 0.949 1.044 0.957 0.000 1.042 1.088 1.092 1.090 0.955 1.031
FT63 0.893 0.946 0.000 0.978 1.009 0.976 0.898 0.988 0.919 0.960 0.000 1.044 1.048 1.046 0.917 0.989
G-EDF 0.950 1.008 1.066 0.000 1.030 0.971** 0.955 1.052 0.929 0.970 1.011 0.000 1.003 1.002 0.927 0.999
G-N 0.923 0.979 1.036 0.971*** 0.000 0.969*** 0.928 1.022 0.926 0.968 1.008 0.997*** 0.000 0.998*** 0.925 0.997
G-t 0.952 1.011 1.069 1.002 1.033 0.000 0.957 1.055 0.928 0.969 1.010 0.999*** 1.002 0.000 0.926 0.998
GAS 0.997 1.058 1.120 1.094 1.128 1.091 0.000 1.105 1.003 1.048 1.093 1.142 1.146 1.144 0.000 1.080
Historical 0.905 0.959 1.014 0.991 1.021 0.988 0.909 0.000 0.929* 0.971 1.012 1.055 1.059 1.057 0.927* 0.000
VaR (5%) VaR (10%)
FT1 0.000 1.034 1.068 1.167 1.167 1.170 0.993 1.060 0.000 1.021 1.046 1.188 1.193** 1.191 0.992* 1.037 1.009 1.029 1.043
FT21 0.967 0.000 1.033 1.128 1.128* 1.130 0.960 1.025 0.980 0.000 1.025 1.163 1.167** 1.165 0.971* 1.016 0.988 1.008 1.022
FT63 0.937 0.968 0.000 1.092 1.091 1.094 0.930 0.993 0.956 0.976 0.000 1.135 1.139 1.137 0.948 0.991 0.965 0.984 0.997
G-EDF 0.915 0.946 0.977 0.000 1.000*** 1.001 0.908 0.969 0.908 0.927 0.949 0.000 1.005*** 1.001 0.900 0.940 0.916 0.934 0.946
G-N 0.914 0.945 0.976 0.999 0.000 1.001 0.908 0.968 0.902 0.921 0.943 0.995 0.000 0.996 0.894 0.934 0.910 0.928 0.940
G-t 0.914 0.945 0.976 0.999*** 0.999*** 0.000 0.908 0.968 0.907 0.926 0.949 0.999*** 1.004*** 0.000 0.899* 0.940 0.915 0.933 0.946
GAS 1.008 1.042 1.076 1.177 1.176 1.179 0.000 1.068 1.009 1.030 1.055 1.198 1.203** 1.200 0.000 1.046 0.973 0.992 1.006
Historical 0.944*** 0.976*** 1.008** 1.099** 1.098** 1.101* 0.937*** 0.000 0.965*** 0.985*** 1.009*** 1.144*** 1.148*** 1.146*** 0.957*** 0.000 0.973* 0.992 1.006
@@ 6 Conclusion and discussion
The question of whether pre-trained foundation time-series models can outperform traditional econometric methodsis fundamentally important from the perspective of economics literature. In this paper, we evaluated the performance of Google’s foundation model for time series, TimesFM, in estimating Value-at-Risk (VaR) for the S&P 100 index and its 91 constituents. We initially used the pre-trained TimesFM model without any task-specific training as a zero-shot predictor and observed significant statistical improvements after fine-tuning it with our data. Our main finding is that in comparison to the state-of-the-art econometric models for VaR, fine-tuned TimesFM performs surprisingly well. It outperforms all econometric benchmarks, including the GAS model, which is widely regarded as the leading method in recent econometric literature (Patton et al., 2019), in terms of Actual vs. Expected violations (AE). Traditional models like GARCH (with different residual distributions) and empirical quantile estimates fall notably behind. These results remain consistent across different VaR levels, underscoring the robustness of TimesFM. Additionally, we assessed performance using the Quantile Score Loss Function, where TimesFM showed results comparable to the best econometric models.
@@
As the results favor the foundation AI model, financial time-series modeling could shift toward data-driven approaches, partially reducing the need for mathematical modeling. Such a shift would have significant implications for both academia and the financial industry, as the application of foundation models can be applied with relatively minimal methodological expertise in econometrics. However, the adoption of foundation models is accompanied by notable challenges. A key challenge with these models is their “black-box” nature, which can make it difficult to understand how they arrive at their predictions. This is particularly concerning because we still have only a partial understanding of how these models operate, where they might fail, and what they are truly capable of, mainly due to their complex, emergent behaviors. While time series foundation models show promise in adapting to out-of-distribution scenarios, similar to how humans use “common sense” reasoning, they are also prone to hallucinations and may produce decisions that appear reasonable but are incorrect (Chakraborty et al., 2024). This lack of transparency is especially problematic in high-stakes financial settings, where it is crucial to understand the reasoning behind predictions for effective risk management and decision-making. Moreover, it can create hurdles for regulatory acceptance, as regulators typically require clear and explainable models to meet compliance standards. Therefore, improving the interpretability of foundation models is critical to align with regulatory expectations and support their wider use in practice.
@@
Overall, our results—which are highly generalizable, having been conducted across over 90 assets with multiple settings—favor the pre-trained foundation model, highlighting its potential to enhance Value-at-Risk (VaR) estimation by offering a completely new alternative to traditional econometric methods. We anticipate further research expanding into various econometric and financial challenges, such as portfolio optimization and volatility forecasting, but studies that examine the risks associated with the use of foundation models on econometric problems closely and critically.

### v3

@@ 4 Evaluation metrics
where the expected violations are calculated as $E=T\times\alpha$, with $T$ being the total number of observations and $\alpha$ the confidence level of the VaR model. Clearly, an AE ratio of 1 is desirable, as it indicates that the model is accurately calibrated. On the other hand, a value $>1$ ($<1$) suggests the underestimation (overestimation) of the risk, both of which are problematic in the sense that an underestimate leads to unexpected negative returns, while the latter leads to high capital requirements, thus, both implying capital losses. Alternatively, a model is deemed good if $|1-\text{AE}|$ is close to zero.
An alternative to AE ratio is the proportion of failure test (POF), also known as the unconditional coverage test (UC). which statistically evaluates whether the actual number of VaR violations matches the expected number of violations over time, based on the chosen confidence level. For instance, it calculates the failure rate defined as the proportion of number of violations ($A$) to the total number of observation $T$ and examines its consistency with the models confidence level $\alpha$. More specifically, let $I_{t}$ denotes the indicator function that the realized return was less than the VaR estimate, i.e.,
@@
In the evaluation of VaR forecasts, the quantile score is a commonly used metric and is now standard in the VaR literature (Ehm et al. 2016, Gneiting 2011, Taylor 2019). The above mentioned statistical tests are good to evaluate the statistical adequacy of VaR estimates, that is, a model for VaR forecasting is deemed to be an adequate model if we fail to reject the null hypothesis. However, these tests cannot conclude whether an ‘adequate’ model outperforms another ‘adequate’ model. Scoring functions allow us to do an honest assessment of VaR as a perfect VaR estimate would always minimize the scoring function. The quantile score $QS_{t}$ can be defined as follows:
@@
For every instant $t$, a model’s score increases either by (a) $\alpha-1$ fraction of the distance between the realized return and the VaR forecast, if realized return is smaller than the VaR forecast for the same period, or (b) by $\alpha$ fraction of the distance otherwise. This scoring rule is consistent for quantile estimation, meaning that it achieves its minimum value when the quantile forecast $q_{t}$ correctly estimates the true quantile of the distribution. Therefore, models that produce lower average quantile scores in the out-of-sample period are considered superior. Finally, to compare the performance of two models, we use the Diebold–Mariano (DM) test (Diebold and Mariano 2002) with
@@ 5.1 Actual over Expected Ratio
Table 2 reports the summary statistics, across 92 stocks, of the $|1-\text{AE}|$ values for the VaR forecasts at different levels, and the corresponding models in the out-sample period. Specifically, we report the minimum, mean, median, maximum, and standard deviation across all the stocks. Additionally, we report the count of stocks for which each of the considered models was the best (achieved lowest value of $|1-AE|$) or was within top two models (1st-2nd best). Note that, as mentioned before, the VaR forecasts for TimesFM pretrained model are only available at $10\%$, there we have three extra models in the last block of Table 2.
@@
where $i$ denotes the model in the selected row, whereas $j$ denotes the model in the selected column. Further, the formatting in the Tables are as follows: *, ** and *** denote whether the $t$-test of equal predictive accuracy is rejected at the 5%, 2.5% and 1% level of significance. Table 3 reports the results to $t$-tests. We can observe that the fine-tuned models, in particular, FT1 and FT21, perform remarkably well across all quantile levels, depicted by failure to reject the null hypothesis in the first and second rows of each block. Further, we observe that FT1 outperforms GAS model in forecasting VaR(1%), VaR(5%) and VaR(10%) at 97.5%, 95% and 99% level of significance respectively, and performs atleast as good as GAS model for forecasting Var(2.5%) (refer rows corresponding to GAS model). In addition, we observe that selecting a short horizon as prediction length performs better in predicting VaR. What’s more, the performance of pre-trained models is not so exciting compared to their fine-tuned counterparts. These findings show the potential of time-series foundation models in forecasting value at risk.
Table 2: Summary statistics of the $|1-AE|$ values over the out-of-sample period from January 2015 to September 2023 for the eleven models. Additionally, we report the count of stocks for which each of the considered models was the best (achieved lowest value of $|1-AE|$) or was within the top two models (1st-2nd best). In case of a tie, equal ranks were given. The values are highlighted using bold for the best values and italicized for the second-best in each column.
@@
Tables 4, 5 and 6 present the results of these tests, namely UC, CC and DQ respectively, for various confidence levels. The results are also depicted via bar plots in Figures 1-3. For each $VaR_{\alpha}$, we present for how many assets we failed to reject the null hypothesis of correct model performance at various confidence levels. The higher the count better the model. From these tables, it is evident that the TimesFM shows promising results; in particular, the fine-tuned TimesFM model with one prediction length (FT1) either comfortably beats or performs fairly in comparison to the benchmark models across all the three tests. More specifically, for the UC test, the fine-tuned TimesFM models with various prediction lengths and the historical model showed the highest number of assets where the null hypothesis was not rejected, indicating that these models had the most consistent unconditional coverage. What’s more remarkable is that the simple rolling method (historical) outperforms the GAS model in UC test. Again, in conditional coverage tests, the FT model with 1 prediction length performed significantly better than all other models. Finally, the performance of FT1 in DQ test is comparable to the benchmark models if not better. In conclusion, the FT1 model demonstrates strong performance, providing robust competition to the GAS model across multiple tests, while TimesFM models with longer prediction lengths show significant weaknesses, particularly in capturing the conditional and dynamic aspects of the VaR process.
@@ 5.2 Quantile Loss
Table 7 presents the summary statistics of quantile loss scores across 92 stocks. We observe that, at 1% VaR, the mean and median quantile score is consistent across the models, with FT1 and FT21 exhibiting lower mean values, 0.062 and 0.065, respectively. This indicates a better performance of fine-tuned TimesFM models compared to GARCH models (G-EDF, G-t, G-N) that have higher mean values and higher values for maximum quantile score. Further, a higher value of the standard deviation for GARCH models shows that these models have high variability in their VaR forecasts which, in turn, reflects that these models are less stable in extreme market scenarios. These observations remain valid at 2.5% VaR forecasts, with FT1 and FT21 showing lower mean and median quantile scores, while GARCH models still show higher values for maximum quantile score and the standard deviation. For VaR at 5% and 10% levels, the historical model performs similar to fine-tuned TimesFM model whereas GARCH models continue to show higher maximum quantile score and greater variability, thus making them less robust and more sensitive to extreme market conditions. In addition, GAS models obtain lower values of mean and median quantile scores, similar to those obtained by FT models across all quantile levels. FT models are thus competitive to the benchmark models, in particular the GAS model which shows slightly higher values of maximum quantile loss and higher standard deviation, particularly at VaR 5% and 10% levels.
@@
As the quantile scores are not easy to interpret, we calculate skill scores defined as the ratio of each model’s score to the score of a benchmark model for an individual asset, and then, we take the mean across assets to summarize the performance of models across multiple assets. More specifically, we report in Table 8, the out-of-sample total quantile loss of each model relative to the benchmark model in the selected row. For instance, for VaR (1%), the first row contains the cross-sectional average of the ratio of out-of-sample total quantile loss of every model relative to the out-of-sample total quantile loss of model FT21. In addition to the comparison of the total quantile loss, we also perform the Diebold–Mariano test for each pair of models where we test
@@
where $i$ denotes the model in the selected row, whereas $j$ denotes the model in the selected column. Further, the formatting in the Tables is as follows: *, **, and *** denote whether the Diebold–Mariano (DM) test of equal predictive accuracy is rejected for more than 50% of the assets at the 5%, 2.5% and 1% level of significance.
VaR (1%)
FT1 FT21 FT63 G-EDF G-N G-t GAS Historical PT1 PT21 PT63
FT1 0.000 1.061 1.123 1.097 1.131 1.095 1.005 1.109
FT21 0.944 0.000 1.058 1.034 1.066 1.032 0.949 1.044
FT63 0.893 0.946 0.000 0.978 1.009 0.976 0.898 0.988
G-EDF 0.950 1.008 1.066 0.000 1.030 0.971** 0.955 1.052
G-N 0.923 0.979 1.036 0.971*** 0.000 0.969*** 0.928 1.022
G-t 0.952 1.011 1.069 1.002 1.033 0.000 0.957 1.055
GAS 0.997 1.058 1.120 1.094 1.128 1.091 0.000 1.105
Historical 0.905 0.959 1.014 0.991 1.021 0.988 0.909 0.000
VaR (2.5%)
FT1 0.000 1.045 1.090 1.138 1.142 1.141 0.998 1.078
FT21 0.957 0.000 1.042 1.088 1.092 1.090 0.955 1.031
FT63 0.919 0.960 0.000 1.044 1.048 1.046 0.917 0.989
G-EDF 0.929 0.970 1.011 0.000 1.003 1.002 0.927 0.999
G-N 0.926 0.968 1.008 0.997*** 0.000 0.998*** 0.925 0.997
G-t 0.928 0.969 1.010 0.999*** 1.002 0.000 0.926 0.998
GAS 1.003 1.048 1.093 1.142 1.146 1.144 0.000 1.080
Historical 0.929* 0.971 1.012 1.055 1.059 1.057 0.927* 0.000
VaR (5%)
FT1 0.000 1.034 1.068 1.167 1.167 1.170 0.993 1.060
FT21 0.967 0.000 1.033 1.128 1.128* 1.130 0.960 1.025
FT63 0.937 0.968 0.000 1.092 1.091 1.094 0.930 0.993
G-EDF 0.915 0.946 0.977 0.000 1.000*** 1.001 0.908 0.969
G-N 0.914 0.945 0.976 0.999 0.000 1.001 0.908 0.968
G-t 0.914 0.945 0.976 0.999*** 0.999*** 0.000 0.908 0.968
GAS 1.008 1.042 1.076 1.177 1.176 1.179 0.000 1.068
Historical 0.944*** 0.976*** 1.008** 1.099** 1.098** 1.101* 0.937*** 0.000
VaR (10%)
FT1 0.000 1.021 1.046 1.188 1.193** 1.191 0.992* 1.037 1.009 1.029 1.043
FT21 0.980 0.000 1.025 1.163 1.167** 1.165 0.971* 1.016 0.988 1.008 1.022
FT63 0.956 0.976 0.000 1.135 1.139 1.137 0.948 0.991 0.965 0.984 0.997
G-EDF 0.908 0.927 0.949 0.000 1.005*** 1.001 0.900 0.940 0.916 0.934 0.946
G-N 0.902 0.921 0.943 0.995 0.000 0.996 0.894 0.934 0.910 0.928 0.940
G-t 0.907 0.926 0.949 0.999*** 1.004*** 0.000 0.899* 0.940 0.915 0.933 0.946
GAS 1.009 1.030 1.055 1.198 1.203** 1.200 0.000 1.046 0.973 0.992 1.006
Historical 0.965*** 0.985*** 1.009*** 1.144*** 1.148*** 1.146*** 0.957*** 0.000 0.973* 0.992 1.006
@@ 6 Conclusion and discussion
Determining whether pre-trained foundation time-series models can outperform traditional econometric methods is a key question, as their success would mean that anyone with minimal statistical and mathematical knowledge could achieve relatively good VaR estimates. In this paper, we evaluated the performance of Google’s foundation model for time series, TimesFM, in estimating Value-at-Risk (VaR) for the S&P 100 index and its 91 constituents. We initially used the pre-trained TimesFM model without any task-specific training that yielded suboptimal results. However, fine-tuning the model with domain-specific data led to significant statistical improvements: In terms of Actual vs. Expected violations (AE), fine-tuned TimesFM outperforms all econometric benchmarks, including the GAS model, which is widely regarded as the leading method in the recent econometric literature (Patton et al. 2019). Traditional models like GARCH (with different residual distributions) and empirical quantile estimates fall notably behind. These results remain consistent across different VaR levels, underscoring the robustness of TimesFM. Additionally, we assessed performance using the Quantile Score Loss Function, where TimesFM showed results comparable to the best econometric models.
@@
As the results favor the foundation AI model, financial time-series modeling could shift toward data-driven approaches, partially reducing the need for mathematical modeling. Such a shift would have significant implications for both academia and the financial industry, as foundation models can be utilized with relatively minimal methodological expertise in econometrics. However, the adoption of foundation models is accompanied by notable challenges. A key challenge with these models is their “black-box” nature, which can make it difficult to understand how they arrive at their predictions. This is particularly concerning because we still have only a partial understanding of how these models operate, where they might fail, and what they are truly capable of, mainly due to their complex, emergent behaviors. While time series foundation models show promise in adapting to out-of-distribution scenarios, similar to how humans use “common sense” reasoning, they are also prone to hallucinations and may produce decisions that appear reasonable but are incorrect (Chakraborty et al. 2024). This lack of transparency is especially problematic in high-stakes financial settings, where understanding the reasoning behind predictions is essential for effective risk management and decision-making. Moreover, it can create hurdles for regulatory acceptance, as regulators typically require clear and explainable models to meet compliance standards. Therefore, improving the interpretability of foundation models is critical to align with regulatory expectations and support their wider use in practice.
@@
Overall, our results—which are highly generalizable, having been conducted across over 90 assets with multiple settings—favor the pre-trained foundation model, highlighting its potential to enhance Value-at-Risk (VaR) estimation by offering a completely new alternative to traditional econometric methods. We anticipate future research addressing various econometric and financial challenges, such as portfolio optimization and volatility forecasting, while also critically examining the risks associated with applying foundation models to econometric problems.
{APPENDICES}
Figure 1: This figure shows the bar plots of number of assets (out of 92) for which we failed to reject the null hypothesis in the Unconditional Coverage (UC) test at various Value at Risk (VaR) levels. A higher number indicates better model performance, as it reflects a closer alignment between observed and expected VaR violations for each specified model. Results are reported for four VaR confidence levels (1%, 2.5%, 5%, and 10%) across different models.
Figure 2: This figure shows the bar plot of the number of assets (out of 92) for which we failed to reject the null hypothesis in the Conditional Coverage (CC) test at various Value at Risk (VaR) levels. A higher number indicates better model performance, as it reflects a closer alignment between observed and expected VaR violations for each specified model. Results are reported for four VaR confidence levels (1%, 2.5%, 5%, and 10%) across different models.
Figure 3: This figure shows the bar plot of the number of assets (out of 92) for which we failed to reject the null hypothesis in the Dynamic Quantile (DQ) test at various Value at Risk (VaR) levels. A higher number indicates better model performance, as it reflects a closer alignment between observed and expected VaR violations for each specified model. Results are reported for four VaR confidence levels (1%, 2.5%, 5%, and 10%) across different models.